Alteran Corporation purchased office equipment for $1.7 million in 2015. The equipment is being depreciated over a 8-year life using the sum-of-the-years'-digits method. The residual value is expected to be $800,000. At the beginning of 2018, Alteran decided to change to the straight-line depreciation method for this equipment.

Required:
Prepare the journal entry to record depreciation for 2018. (If no entry is required for a transaction/event, select "No journal entry required" in the first account field. Enter your answers in whole dollars.)
Answer #1

Accumulated depreciation = (1700000-800000)*21/36 = 525000

2018 Depreciation = (1700000-525000-800000)/5 = 75000

Date account and explanation Debit Credit
2018 Depreciation expense 75000
Accumulated depreciation-equipment 75000